Cesar is 101 years old. Taylor is 41 years old. How many years ago was Cesar's age 4 times Taylor's age?
Found 2 solutions by Alan3354, fractalier:
Answer by Alan3354(69443)

Cesar is 101 years old. Taylor is 41 years old. How many years ago was Cesar's age 4 times Taylor's age?
===============
The age difference is 60 years.
--> When they were 20 & 80

Answer by fractalier(6550)

Call the number of years ago you want, x. Then we write
101 - x = 4(41 - x)
101 - x = 164 - 4x
3x = 63
x = 21 years ago
They were 80 and 20.